OpenStudy (anonymous):

if radical n is equal to 9, what is the value of n

sam (.sam.):

\[\sqrt{n}=9 \\ \\ (\sqrt{n})^2=9^2 \\ \\ n=?\]

OpenStudy (anonymous):

Thanks Sam

sam (.sam.):

yw